# Electronic correlation effects in the response of graphene and MoS2 monolayers to the impact of highly-charged ions

### v1

@@ 2.4.1 Embedding Hamiltonian
The Hamiltonian of the total system can be written in the following form [31]
@@ 2.4.2 NEGF embedding equations
where we drop the superscripts, in the following. This ansatz enables us to write the equations of motion for the whole system as [33]
@@
The embedding approach allows one to derive a closed equation for the system Green function, Eq. (9), where the e-s coupling is exactly accounted for via an additional “embedding” self-energy, Eq. (10), see also Ref. [17]. However, this equation still has the unfavorable cubic scaling of the CPU-time with $N_{t}$ that is inherent in the two-time KBE.
Therefore, it is again advantageous to derive a time-local G1–G2 version of the embedding scheme, as was demonstrated in Ref. [33]. The main steps are again: transition to the real-time plane, application of the HF-GKBA and differentiation of the resulting expressions with respect to time. The result is the following coupled set of time-local EOMs
@@
For details of the derivation and conservation properties, we refer to Ref. [33].
@@ 2.5 Derivation of the embedding equations in a time-local framework
In the following, we use the above equations in the G1–G2 embedding framework and apply them to the ion stopping scenario that was discussed before.
@@ 3.1 Model Hamiltonians for graphene and MoS2 monolayers
In Sec. 2, we introduced the general form of the many-body Hamiltonian. To implement a specific model, we now have to find expressions for the single-particle Hamiltonian, as well as the pair potential. Ultimately we want to describe the electron dynamics in 2D graphene and MoS2 monolayers. For graphene we use the approach that was introduced in Refs. [32, 6], where the delocalised $\pi$ electrons are treated by a single-band Hubbard model
@@
where also an external potential $W_{ii}$ is included. This model depends solely on three physical parameters: the on-site energy $\epsilon$, the nearest-neighbor hopping amplitude $J$ and the Hubbard interaction strength $U$. The geometry of the lattice is captured by the nearest-neighbor sum, indicated by $\braket{i,j}$ and the full, pairwise Coulomb interaction is approximated by the local, on-site, repulsive energy contribution proportional to the spin densities $\hat{n}_{i,\sigma}$. In this work we use the numerical values $\epsilon=1.22J_{0}$, $J=0.74J_{0}$, and $U=1.19J_{0}$ [6], which are obtained by fitting the resulting band structure to DFT results where we adopt the unit system $J_{0}=\hbar/m_{\mathrm{e}}/a_{0}^{2}\approx 3.78\mathrm{eV}$, $a_{0}\approx 1.42$Å [32].
Now we discuss the case of MoS2 in more detail. While this single-band approach is suitable for single-layer graphene (SLG), MoS2 has a more complex band structure. Efforts to reduce the full band structure into compressed 5-band [41] and 3-band [42] models have proven successful in reproducing the DFT band structure. However, these models are tight-binding models, i.e. they completely neglect correlations. The over-fitted band structure does not leave room for a simple interaction model, which is necessary to study the effects of correlations in the G1–G2 scheme.
When projected onto the plane containing the monolayer, the S-Mo-S unit cells form a 2D honeycomb lattice. We, therefore, employ the same strategy as in Ref. [6] and use the same effective, single-band Hamiltonian to describe four identical, non-interacting electron bands in MoS2. The parameters which provide a good fit for the physical system are $\epsilon=1.19J_{0}$, $J=0.29J_{0}$ and $U=1.19J_{0}$. The lower ratio of $J/U$ in MoS2, compared to graphene, reflects the low carrier mobility in the semiconductor.
@@
In the absence of interactions ($U=0$) the density of states (DOS) of the system can be obtained directly by diagonalization of the Hamiltonian. An example of honeycomb flakes of different size (colors) are shown in Fig. 1 a). Finite size effects lead to so called edgestates, which appear as secondary peaks in the DOS (Fig. 1, b)). Their relative magnitude decreases with the system size and the shape of the DOS approaches the TB-DOS for the infinite lattice, also depicted in Fig. 1, b) for comparison.
@@ 3.2 Model Hamiltonian for the ion stopping and charge transfer
In the presence of a highly-charged ion (HCI) with initial charge $Z_{0}$ we use a semiclassical Ehrenfest model according to which the Hamiltonian is modified by an additional diagonal contribution, due to the Coulomb potential of the ion [29, 32]
$W_{ii}=-\frac{Z_{0}W_{0}a_{0}}{|\bm{r}-\bm{S}(t)|},$ (15)
with the rescaled Coulomb constant $W_{0}=e^{2}/(4\pi\epsilon_{0}a_{0})\approx 2.68J_{0}$. In this paper we focus on simple straight ion trajectories, normal to the cluster plane, $\bm{S}(t)=(z_{0}+v_{0}t)\bm{e}_{z}$, without updating the velocity, $v_{0}$, which is well justified for the comparatively high ion velocities studied below. To reduce the complexity of the model, we neglect off-diagonal contributions. The HCI itself will be described by a set of static energy levels
@@
derived from the value of the Coulomb potential at the point of resonant charge transfer. This implies that electrons on the ion can not decay into lower orbitals, eliminating complex processes, such as Auger recombination or interatomic Coulomb decay [43] from this picture.
The charge transfer (ion neutralization) Hamiltonian will be described by a Gaussian amplitude for the transition of electrons from specific system sites into the environment sites (i.e., ionic orbitals) [32]
$\hat{H}^{\mathrm{se}}=\gamma[\bm{S}(t)]\sum_{i_{1},i_{2},\sigma}\delta^{\mathrm{CT}}_{i_{1},i_{2}}\hat{c}_{i_{1},\sigma}^{\dagger}\hat{a}_{i_{2},\sigma}\,,\quad\gamma[\bm{S}(t)]=\gamma_{0}\exp\left(\frac{(z(t)+z_{\mathrm{res}})^{2}}{2d_{\mathrm{w}}^{2}}\right).$ (17)
The peak of the Gaussian is located at the point of resonant charge transfer, $z_{\mathrm{res}}=-\sqrt{3}a_{0}$, has a width of $d_{\mathrm{w}}=0.6a_{0}$ and an amplitude of $\gamma_{0}=2.12$, that was determined in Ref. [32]. The charge transfer Kronecker delta, $\delta^{\rm CT}$, determines the coupling scheme between the system and the HCI. In order to mimic realistic charge transfer, the sites nearest to the HCI penetration point are coupled to the ion, and since each site can contribute 2 electrons per each of the four bands, the number of coupled sites is chosen to be $Z_{0}/8$ (integer). This allows for full neutralization of the ion, while limiting the charge flux (see Fig. 2).
@@
Figure 3: Total charge transfer as a function of the inverse ion velocity for different charge states. a) SLG: comparison between NEGF [33] and the present time-local embedding scheme, both in HF approximation, as well as experimental data [4]. b) Comparison between the present simulations for SLG and MoS2, compared to experimental data for MoS2 [44].
@@
Let us now compare the charge transferred to the ion between the two materials, single-layer graphene (SLG) and MoS2, cf. Fig. 3, b. Our results are depicted by the colored lines with open circles (SLG) and crosses (MoS2) and predict a systematically slightly higher charge transfer for the case of MoS2 which is in agreement with the experimental observations of Ref. [44]. There it was noted that, an important reason for the increased charge transfer is the larger layer width of MoS2 leading to a larger interaction time with the ion. In our model no finite monolayer width effects are included and the increased charge transfer of MoS2 results from its larger lattice constant compared to SLG. This will be discussed in more detail below, in the context of Fig. 4. Compared to the experiments our data for the charge transfer appear to be slightly higher, so we also repeated simulations with a slightly reduced charge transfer amplitude $\gamma_{0}$ compared to the one determined in Ref. [32], however, this did not have a noticable effect on the electron dynamics in the monolayer. Therefore, in all simulations of this paper we used the same amplitude $\gamma_{0}=2.12J_{0}$, for both materials.
@@ 4.1 Ion-induced electron dynamics in the target
where $R$ is the set of $L_{R}$ sites which compose a ring. During the approach of the HCI, the central rings accumulate negative charge due to the attractive Coulomb potential of the ion. Subsequently, the outer rings get depleted, resulting in a positive charge. The inclusion of correlations has a significant effect on the charge distribution in the MoS2 cluster: the charge transfer to the ion is reduced and electron migration between the rings is smoothened, see Fig. 4, a). At the point of resonant charge transfer, electrons are transferred from the cluster onto the HCI. This transfer is very rapid (sub-femtosecond) in both approximations. During the interaction time with the HCI, a period of constant, positive charge in the center of the cluster can be observed, during which the HCI partially neutralizes. Once the interaction with the ion is over (positive times), electrons flow back into the center of the cluster from outer layers as a plasmon-like wave. Interestingly, the onset of this flow is shifted to later times (more than 1fs) in the HF approximation compared to SOA. Also note the high-frequency oscillations in the HF approximation, which are dampened out by the inclusion of correlations.
@@ 4.3 Ion-induced doublon dynamics in the monolayer
This quantity is important for the charge transport in correlated Fermi systems. For example, after a confinement quench in fermionic atoms doublons and single electrons have been found to diffuse with different speeds [1, 2]. It is interesting to spatially resolve the doublon number. On a hexagonal ring with index $R$ their number is given by
@@
Driving a strongly-correlated system out of equilibrium can lead to the formation of a significant number of doublons, which was observed for the impact of a weakly charged ion or multiple ions [30, 45]. An even stronger effect is observed in the present excitation scenario with highly-charged ions: during the approach of the HCI the doublon occupation in the central rings increases sharply because of the attractive Coulomb potential of the ion that acts like a time-dependent confinement in the monolayer. Unlike in Ref. [30], this increased occupation does not decay monotonically towards a stable configuration. This is because electrons from the central ring are transferred to the HCI which leads to a rapid decrease of the doublon number in the vicinity of the ion impact point. After the interaction time between the HCI and the cluster (after about 2.5 fs), the doublon occupation on the innermost ring increases again and subsequently slowly decays towards the initial value before the ion impact (see Fig. 6). In contrast, in HF, this relaxation is much more violent and oscillatory representing artificial effects due to the neglect of correlations.
@@ 4.4 Induced electrostatic potential
An important quantity that was found to crucially influence the emission of secondary electrons after ion impact is the electrostatic potential induced in the cluster as a consequence of the ion Coulomb force and the resonant charge transfer [6]. In this reference it was observed that this potential is primarily responsible for the dramatic differences in the number of secondary electrons (more than a factor 8 higher in SLG) emitted from SLG and MoS2, respectively. The space and time-dependent potential follows from the solution of Poisson’s equation containing the time-dependent charge density $\rho_{i}(t)$ in the cluster, where $i$ labels the lattice sites,
@@
A snapshot of the induced potential at time t=2t=2fs and at a height z=0.25z=0.25Å above the monolayer is shown in Figs.  7.b)-e), where the corresponding instantaneous charge densities can be read-off from Fig. 4, cf. the vertical dotted line. These parameters are chosen because they allow us to directly compare to the HF results that were presented in Fig. 2 of Ref. [6]. First, a strongly non-monotonic behavior in radial direction $x$ is evident where the extrema correspond to the mean shell radii. The innermost peak is located at the shell from which electrons are transferred to the ion. Notice the striking differences in figure parts b) and c) corresponding to SLG and MoS2, respectively: For SLG this innermost peak is a local minimum whereas, for MoS2, it is a local maximum. This has dramatic consequences for secondary electron emission [6]: Electrons that are generated via the interatomic Coulombic decay (ICD) are subject to the potential $V$ and the associated force on secondary electrons (SE), $\textbf{F}^{\rm SE}$, Eq. (23). Due to the $r^{-6}$ dependence on the distance between ion and emitting lattice site, ICD generation of secondary electrons will only be effective from the innermost ring. Thus, an ICD generated electron will be able to leave the monolayer towards the detector if the $z$-component of the force is positive (repulsive). Since the potential $V$ overall, decays with increasing height $z$, the $z$-component of the force is typically positive. However, close to the monolayer deviations from this trend occur: While, in case of SLG, the local minimum at the inner ring increases the repulsive force, the local maximum, in case of MoS2, in contrast, gives rise to an attractive force, in the vicinity of the monolayer, see Fig. 8. This explains the much larger number of secondary electrons that are emitted from SLG as compared to MoS2 [6].

### v2

@@ 2.4.1 Embedding Hamiltonian
The Hamiltonian of the total system can be written in the following form [Bonitz_fcse_19]
@@ 2.4.2 NEGF embedding equations
where we drop the superscripts, in the following. This ansatz enables us to write the equations of motion for the whole system as [balzer_prb_23]
@@
The embedding approach allows one to derive a closed equation for the system Green function, Eq. (10), where the e-s coupling is exactly accounted for via an additional “embedding” self-energy, Eq. (11), see also Refs. [myohanen_prb_09, stefanucci_many-body_2013]. However, this equation still has the unfavorable cubic scaling of the CPU-time with $N_{t}$ that is inherent in the two-time KBE.
Therefore, it is again advantageous to derive a time-local G1–G2 version of the embedding scheme, as was demonstrated in Ref. [balzer_prb_23]. The main steps are again: transition to the real-time plane, application of the HF-GKBA and differentiation of the resulting expressions with respect to time. The result is the following coupled set of time-local EOMs
@@
For details of the derivation and conservation properties, we refer to Ref. [balzer_prb_23]; a similar result has been obtained in Ref. [tuovinen_prl_23].
@@ 2.5 Derivation of the embedding equations in a time-local framework
In the following, we use the above equations in the G1–G2 embedding framework and apply them to the ion stopping scenario that was discussed before. We will concentrate on ultrafast charge transfer processes that proceed within a few femtoseconds. On these time scales the lattice response is only of minor importance (the phonon period of MoS2 is on the order of 80 fs) and will not be considered here. Phonon processes can be directly incorporated in our time linear scheme as was demonstrated e.g. in Refs. [stefanucci_prx_23, emeis_prx_25].
@@ 3.1 Model Hamiltonians for graphene and MoS2 monolayers
In Sec. 2, we introduced the general form of the many-body Hamiltonian. To implement a specific model, we now have to find expressions for the single-particle Hamiltonian, as well as the pair potential. Ultimately we want to describe the electron dynamics in 2D graphene and MoS2 monolayers. For graphene we use the approach that was introduced in Refs. [balzer_cpp_21, niggas_prl_22], where the delocalised $\pi$ electrons are treated by a single-band fourfold degenerate Hubbard model
@@
where also an external potential $W_{ii}$ is included. This model depends solely on three physical parameters: the on-site energy $\epsilon$, the nearest-neighbor hopping amplitude $J$ and the Hubbard interaction strength $U$. The geometry of the lattice is captured by the nearest-neighbor sum, indicated by $\braket{i,j}$ and the full, pairwise Coulomb interaction is approximated by the local, on-site, repulsive energy contribution proportional to the spin densities $\hat{n}_{i,\sigma}$. In this work we use the numerical values $\epsilon=1.22J_{0}$, $J=0.74J_{0}$, and $U=1.19J_{0}$ [niggas_prl_22], which are obtained by fitting the resulting band structure to DFT results where we adopt the unit system $J_{0}=\hbar/m_{\mathrm{e}}/a_{0}^{2}\approx 3.78\mathrm{eV}$, $a_{0}\approx 1.42$Å [balzer_cpp_21].
Now we discuss the case of MoS2 in more detail. While this single-band approach is suitable for single-layer graphene (SLG), MoS2 has a more complex band structure. Efforts to reduce the full band structure into compressed 5-band [shanavas_effective_2015] and 3-band [liu_three-band_2013] models have proven successful in reproducing the DFT band structure. However, these models are tight-binding models, i.e. they completely neglect correlations. The over-fitted band structure does not leave room for a simple interaction model, which is necessary to study the effects of correlations in the G1–G2 scheme.
When projected onto the plane containing the monolayer, the S-Mo-S unit cells form a 2D honeycomb lattice. We, therefore, employ the same strategy as in Ref. [niggas_prl_22] and use the same effective, single-band Hamiltonian to describe four identical, non-interacting electron bands in MoS2. The parameters which provide a good fit for the physical system are $\epsilon=1.19J_{0}$, $J=0.29J_{0}$ and $U=1.19J_{0}$. The lower ratio of $J/U$ in MoS2, compared to graphene, reflects the low carrier mobility in the semiconductor.
@@
In the absence of interactions ($U=0$) the density of states (DOS) of the system can be obtained directly by diagonalization of the Hamiltonian. An example of honeycomb flakes of different size (colors) are shown in Fig. 1 a). Finite size effects lead to so-called edgestates, which appear as secondary peaks in the DOS (Fig. 1, b)). Their relative magnitude decreases with the system size and the shape of the DOS approaches the TB-DOS for the infinite lattice, also depicted in Fig. 1, b) for comparison.
@@ 3.2 Model Hamiltonian for the ion stopping and charge transfer
In the presence of a highly-charged ion (HCI) with initial charge $Z_{0}$ we use a semiclassical Ehrenfest model according to which the Hamiltonian is modified by an additional diagonal contribution, due to the Coulomb potential of the ion [balzer_prb16, balzer_cpp_21]
$W_{ii}=-\frac{Z_{0}W_{0}a_{0}}{|\bm{r}_{i}-\bm{S}(t)|},$ (16)
with the rescaled Coulomb constant $W_{0}=e^{2}/(4\pi\epsilon_{0}a_{0})\approx 2.68J_{0}$. In this paper we focus on simple straight ion trajectories, normal to the cluster plane, $\bm{S}(t)=z(t)\bm{e}_{z}=(z_{0}+v_{0}t)\bm{e}_{z}$, without updating the velocity, $v_{0}$, which is well justified for the comparatively high ion velocities studied below. To reduce the complexity of the model, we neglect off-diagonal contributions. The HCI itself is modeled by a set of degenerate static energy levels with energy $\epsilon$
@@
derived from the value of the Coulomb potential at the point of resonant charge transfer. This implies that electrons on the ion can not decay into lower orbitals, eliminating complex processes, such as Auger recombination or interatomic Coulomb decay [jahnke_interatomic_2020] from this picture.
The charge transfer (ion neutralization) Hamiltonian will be described by a Gaussian amplitude for the transition of electrons from specific system sites into the environment sites (i.e., ionic orbitals) [balzer_cpp_21]
$\hat{H}^{\mathrm{se}}=\gamma[\bm{S}(t)]\sum_{i_{1},i_{2},\sigma}\delta^{\mathrm{CT}}_{i_{1},i_{2}}\left\{\hat{c}_{i_{1},\sigma}^{\dagger}\hat{a}_{i_{2},\sigma}+\hat{a}^{\dagger}_{i_{2},\sigma}\hat{c}_{i_{1},\sigma}\right\}\,,\quad\gamma[\bm{S}(t)]=\gamma_{0}\exp\left(-\frac{(z(t)+z_{\mathrm{res}})^{2}}{2d_{\mathrm{w}}^{2}}\right).$ (18)
The peak of the Gaussian is located at the point of resonant charge transfer, $z_{\mathrm{res}}=-\sqrt{3}a_{0}$, has a width of $d_{\mathrm{w}}=0.6a_{0}$ and an amplitude of $\gamma_{0}=2.12$, that was determined in Ref. [balzer_cpp_21]. The charge transfer Kronecker delta, $\delta^{\rm CT}$, determines the coupling scheme between the system and the HCI. In order to mimic realistic charge transfer, the sites nearest to the HCI penetration point are coupled to the ion, and since each site can contribute 2 electrons per each of the four bands, the number of coupled sites is chosen to be $Z_{0}/8$ (integer). This allows for full neutralization of the ion, while limiting the charge flux (see Fig. 2).
@@
Figure 3: Total charge transfer as a function of the inverse ion velocity for different charge states. a) SLG: comparison between NEGF [balzer_prb_23] and the present time-local embedding scheme, both in HF approximation, as well as experimental data [gruber_ultrafast_2016]. b) Comparison between the present simulations for SLG and MoS2, compared to experimental data for MoS2 [schwestka_charge_2020].
@@
Let us now compare the charge transferred to the ion between the two materials, single-layer graphene (SLG) and MoS2, cf. Fig. 3, b. Our results are depicted by the colored lines with open circles (SLG) and crosses (MoS2) and predict a systematically slightly higher charge transfer for the case of MoS2 which is in agreement with the experimental observations of Ref. [schwestka_charge_2020]. There it was noted that, an important reason for the increased charge transfer is the larger layer width of MoS2 leading to a larger interaction time with the ion. In our model no finite monolayer width effects are included and the increased charge transfer of MoS2 results from its larger lattice constant compared to SLG. This will be discussed in more detail below, in the context of Fig. 4. Compared to the experiments our data for the charge transfer appear to be slightly higher, so we also repeated simulations with a slightly reduced charge transfer amplitude $\gamma_{0}$ compared to the one determined in Ref. [balzer_cpp_21], however, this did not have a noticable effect on the electron dynamics in the monolayer. Therefore, in all simulations of this paper we used the same amplitude $\gamma_{0}=2.12J_{0}$, for both materials.
@@ 4.1 Ion-induced electron dynamics in the target
where $R$ is the set of $L_{R}$ sites which compose a ring. During the approach of the HCI, the central rings accumulate negative charge due to the attractive Coulomb potential of the ion. Subsequently, the outer rings get depleted, resulting in a positive charge. The inclusion of correlations has a significant effect on the charge distribution in the MoS2 cluster: the charge transfer to the ion is reduced and electron migration between the rings is smoothened, see Fig. 4, a). At the point of resonant charge transfer, electrons are transferred from the cluster onto the HCI. This transfer is very rapid (sub-femtosecond) in both approximations. During the interaction time with the HCI, a period of constant, positive charge in the center of the cluster can be observed, during which the HCI partially neutralizes. Once the interaction with the ion is over (positive times), electrons flow back into the center of the cluster from outer layers as a plasmon-like wave. Interestingly, the onset of this flow is shifted to later times (more than 1fs) in the HF approximation compared to SOA. The reason is that, in SOA, the change of the charge on the rings is significantly reduced compared to HF (except for the second ring). Also note the high-frequency oscillations of the charge density on the rings for $t\gtrsim 2$fs in the HF approximation, which are dampened out by the inclusion of correlations.
@@ 4.3 Ion-induced doublon dynamics in the monolayer
This quantity is important for the charge transport in correlated Fermi systems. For example, after a confinement quench in fermionic atoms doublons and single electrons have been found to diffuse with different speeds [schneider_np_12, schluenzen_prb16]. It is interesting to spatially resolve the doublon number. On a hexagonal ring with index $R$ their number is given by
@@
Driving a strongly-correlated system out of equilibrium can lead to the formation of a significant number of doublons, which was observed for the impact of a weakly charged ion or multiple ions [balzer_prl_18, borkowski_pss_22]. An even stronger effect is observed in the present excitation scenario with highly-charged ions: during the approach of the HCI the doublon occupation in the central rings increases sharply because of the attractive Coulomb potential of the ion that acts like a time-dependent confinement in the monolayer. Unlike in Ref. [balzer_prl_18], this increased occupation does not decay monotonically towards a stable configuration. This is because electrons from the central ring are transferred to the HCI which leads to a rapid decrease of the doublon number in the vicinity of the ion impact point. After the interaction time between the HCI and the cluster (after about 2.5 fs), the doublon occupation on the innermost ring increases again and subsequently slowly decays towards the initial value before the ion impact (see Fig. 6). In contrast, in HF, this relaxation is much more violent and oscillatory representing artificial effects due to the neglect of correlations.
@@ 4.4 Induced electrostatic potential
An important quantity that was found to crucially influence the emission of secondary electrons after ion impact is the electrostatic potential induced in the cluster as a consequence of the ion Coulomb force and the resonant charge transfer [niggas_prl_22]. In this reference it was observed that this potential is primarily responsible for the dramatic differences in the number of secondary electrons (more than a factor 8 higher in SLG) emitted from SLG and MoS2, respectively. The space and time-dependent potential follows from the solution of Poisson’s equation containing the time-dependent charge density $\rho_{i}(t)$ in the cluster, where $i$ labels the lattice sites,
@@
A snapshot of the induced potential at time t=2t=2fs and at a height z=0.25z=0.25Å above the monolayer is shown in Figs.  7.b)-e), where the corresponding instantaneous charge densities can be read-off from Fig. 4, cf. the vertical dotted line. These parameters are chosen because they allow us to directly compare to the HF results that were presented in Fig. 2 of Ref. [niggas_prl_22]. First, a strongly non-monotonic behavior in radial direction $x$ is evident where the extrema correspond to the mean shell radii. The innermost peak is located at the shell from which electrons are transferred to the ion. Notice the striking differences in figure parts b) and c) corresponding to SLG and MoS2, respectively: For SLG this innermost peak is a local minimum whereas, for MoS2, it is a local maximum. This has dramatic consequences for secondary electron emission [niggas_prl_22]: Electrons that are generated via the interatomic Coulombic decay (ICD) are subject to the potential $V$ and the associated force on secondary electrons (SE), $\textbf{F}^{\rm SE}$, Eq. (24). Due to the $r^{-6}$ dependence on the distance between ion and emitting lattice site, ICD generation of secondary electrons will only be effective from the innermost ring. Thus, an ICD generated electron will be able to leave the monolayer towards the detector if the $z$-component of the force is positive (repulsive). Since the potential $V$ overall, decays with increasing height $z$, the $z$-component of the force is typically positive. However, close to the monolayer deviations from this trend occur: While, in case of SLG, the local minimum at the inner ring increases the repulsive force, the local maximum, in case of MoS2, in contrast, gives rise to an attractive force, in the vicinity of the monolayer, see Fig. 8. This explains the much larger number of secondary electrons that are emitted from SLG as compared to MoS2 [niggas_prl_22].